In a population, it was found that the diastolic blood pressure is normally distributed with a mean 81.7 and standard deviation of 7.2 mm Hg. What proportion of the population have diastolic blood pressures above 80 mm Hg?
Answer #1

u=8107, 7.2 P(x > 80 = P(Z >30-4 =HP(7<80 -1852) = 1- P(Z <-8-24) [using z table = 1- 0.4052 0.5948